Required information [The following information applies to the questions displayed below.] Emily Company uses a periodic inveRequired: 1. Prepare a separate income statement through pretax income that details cost of goods sold for (a) Case A: FIFO apls help me fill in the blanks!
Answer #1
Case A Case B
FIFO LIFO
Sales revenue 644280 644280
Cost of goods sold
Beginning inventory 38740 38740
Purchases 272810 272810
Goods available for sale 311550 311550
Ending inventory 161650 119520
Cost of goods sold 149900 192030
Gross Profit 494380 452250
Operating expenses 186500 186500
Pretax income 307880 265750
Workings:
Ending inventory units 8750 =2980+8860+7830-10920
Ending inventory:
FIFO 161650 =(7830*19)+(8750-7830)*14
LIFO 119520 =(2980*13)+(8750-2980)*14